OpenStudy (anonymous):

please help me...... If 6/11 of k is 8/41, what is 3/11 of k?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

lol i meant \[\frac{3}{11}\]is half of \[\frac{6}{11}\] right?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

so what is half of \[\frac{8}{41}\]?

OpenStudy (anonymous):

but isn't that a whole different problem? i dont understand why we have to multiply that by 1/2

OpenStudy (misty1212):

?

OpenStudy (anonymous):

(6/11)k = (8/41) solve for just k and find (3/11) of k

OpenStudy (anonymous):

u multiply 1/2 times 8/41 right?

OpenStudy (misty1212):

yes that is one way to say it

OpenStudy (misty1212):

i would say "take half of it" because it is pretty clear that half of eight is four

OpenStudy (anonymous):

but why ?

OpenStudy (anonymous):

y do you take half of 8/41?

OpenStudy (misty1212):

no need to solve for \(k\) just take half of \(\frac{8}{41}\)

OpenStudy (misty1212):

ok lets to another simpler one

OpenStudy (misty1212):

if \[10\times x=50\] what is \[5\times x\]?

OpenStudy (anonymous):

5?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

if \[10x=50\] then \[x=5\] but if \[10x=50\] then \[5x=25\]

OpenStudy (misty1212):

that is because no matter what \(x\) is, \(5x\) is half of \(10x\)

OpenStudy (anonymous):

o okay

OpenStudy (misty1212):

since \(\frac{3}{11}\) is half of \(\frac{6}{11}\) that means \[\frac{3}{11}k\] is half of \[\frac{6}{11}k\]

OpenStudy (misty1212):

you are told that \[\frac{6}{11}k=\frac{8}{41}\] that makes \[\frac{3}{11}k=\frac{4}{41}\] half of it

OpenStudy (anonymous):

hmmmm..... i think i see now....